Assume a Poisson distribution with-5.8. Find the following probabilities. b.X-1 с. X>1Compute the mean and standard deviation for the following hypergeometric distributions. a. n-3, N 9, and E-7 b. n- 5, N-8, and E-3 c. n-6, N 14, and E 3 d. n 4, N-9, and E 4
Answer #1

(1)

(a)

e-5.35.81 0.0030×5.8 P(X = 1) =-1! 0.0176

(b)

P(X<1) = P(X=0)

5.85.80 0.0030x1-0.0030 P(X = 0) =-O!

So,

P(X<1)= 0.0030

(c)

P(X>1) = 1 - [P(X=0) + P(X=1)]

= 1 - (0.0030 + 0.0176)

= 0.9794

(d)
P(Xleq1) = P(X = 0)+ P(X=1)

= 0.0030 + 0.0176

= 0.0206

(2)

Formula:

p=rac{E}{N}

E(X) = np

n)

(a)

n = 3, N = 9, E = 7

Substituting, we get:

7 p 0.7778

(i)

Mean is given by:

E(X) = np= 3 × 0.7778 = 2.3333

(ii)

Variance is given by:

3 × 0.7778(1-0.7778) (9-3) 9-1 3 × 0.7778 × 0.2222 × 6 Var(X) = 3.1109

So,

Standard Deviation is given by:

V3.1109- 1.7638

(b)

n = 5, N = 8, E = 3

Substituting, we get:

p =-= 0.375

(i)

Mean is given by:

E(X) = np= 5 × 0.375 = 1.875

(ii)

Variance is given by:

5 × 0.375(1-0.375)(8-5) 8-1 5 × 0.375 × 0.625 × 3 7 Var(X) = = 0.5022

So,

Standard Deviation is given by:

V0.5022-0.7087

(c)

n = 6, N = 14, E = 3

Substituting, we get:

p=rac{3}{14}=0.2143

(i)

Mean is given by:

E(X) = np= 6 × 0.2143= 1.2857

(ii)

Variance is given by:

Var(X)=rac{6 imes 0.2143(1-0.2143)(14-6)}{14-1}=rac{6 imes 0.2143 imes 0.7857 imes 8}{13}=0.6217

So,

Standard Deviation is given by:

sqrt{0.6217}=0.7885

(d)

n = 4, N = 9, E = 4

Substituting, we get:

p=rac{4}{9}=0.4444

(i)

Mean is given by:

E(X)=np=4 imes 0.4444=1.7778

(ii)

Variance is given by:

4 × 0.4444(1-0.4444) (9-4) 9-1 4 × 0.4444 × 0.5556 × 5 Var(X) = 0.6173

So,

Standard Deviation is given by:

sqrt{0.6173}=0.7857